The Thakor Corporation’s purchases from suppliers in a quarter are equal to 60 percent of the next quarter’s forecasted sales. The payables period is 60 days. Wages, taxes, and other expenses are 25 percent of sales, and interest and dividends are $95 per quarter. No capital expenditures are planned. Projected quarterly sales are:

Q1 1869

Q2 2160

Q3 1860

Q4 1560

Sales for the first quarter of the following year are projected at $2,190. Calculate the company’s cash outlays by completing the following: (Do not round intermediate calculations and round your answers to 2 decimal places, e.g., 32.16.)

Q1 Q2    Q3 Q4

Payment of accounts $ $ $ $

Wages, taxes, other expenses

Long-term financing expenses (interest and dividends)

Total $ $ $ $
Answer #1

Calculating the Payments on Account Q4 Q2 Q3 Q4 Sales Purchases( 60% of Next Quarter) Paymnets ( 2/3 in Previous uarter,1/3 in Current Quarter 1869 1296 2160 1116 18601560 2190 1314 1121.4 936 1179.6 1236 10561062 Answer Q4 1062.00 390.00 Q2 Q3 Pavements of Accounts Wages, Taxes ,Other Expenses Long-term Financin Total 1179.60 467.25 95.00 1741.85 1236.00 540.00 95.00 1871.00 1056.00 465.00 95.00 1616.00 95.00 1547.00 ensesNote : Payable period is 60 days : Date is provided with the Quarters ( 90 days) :

60 days Means = 2/3+1/3

2 months Payment from Previous quarter Purchases = 2/3 of Previous quarter  Orders

1 month Payment from this Quarter Purchases = 1/3 of Current Quarter Order

Payment on accounts = 2/3 of Previous quarter orders + 1/3 of Current Quarter Order